OpenStudy (anonymous):

The area of a triangle is 92 cm2. The base of the triangle is 8 cm. What is the height of the triangle? A. 11.5 cm B. 23 cm C. 4 cm D. 46 cm help for a medal

OpenStudy (anonymous):

h = 2T / a = 2*92/8 = 23 B is good answer

OpenStudy (anonymous):

T = 1/2 a h (basic rule for area of triangle)